OpenStudy (anonymous):

A box with a square base and open top must have a volume of 4,000 cm3. Find the dimensions of the box that minimize the amount of material used.

OpenStudy (anonymous):

V = 4,000 = x^2 (h), where x is the one side of the base, h is the height and V is the volume S = x^2 + 4(x)(h), where S is the surface area. h = 4,000/(x^2) S = x^2 + 16,000(1/x) differentiate S with respect to x dS/dx = 2x - 16,000/x^2 = 0 since dS/dx = 0 is a minimum 2x^3 - 16,000 = 0 x^3 = 8,000, x = 20 The sides are 20 cm, and the height is 10 cm

OpenStudy (anonymous):

Does that help?

OpenStudy (anonymous):

why did you set the derivative of S = 0

OpenStudy (anonymous):

idk yahoo lol

OpenStudy (anonymous):

hahaha. well thanks anyway lol. the answer came up as correct. I'll try and figure out what was done lol

OpenStudy (anonymous):

lol no problem , good luck with the rest.

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

your welcome!